OpenStudy (anonymous):

76 Ounces=______ Pounds, Round to The Nearest Hundredth..?

OpenStudy (anonymous):

16 ounces=1 pound 76ounces=1/16 *76=76/16 solve it

OpenStudy (anonymous):

1216?

OpenStudy (anonymous):

no it is 19/4=4.75 pounds